OpenStudy (anonymous):

y=2 v1-4v2+v3 find ||y||

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

{v1,v2,v3} is orthonormal basis

OpenStudy (anonymous):

@myininaya

OpenStudy (anonymous):

@samjordon

OpenStudy (anonymous):

isnt magnitude the sqrt of each compenent squared?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

sqrt(2^2+4^2+1^2)=sqrt(21)

OpenStudy (anonymous):

hmmm if it was orthonormal then the magnitude wld be one if i remember correctly

OpenStudy (anonymous):

what about v1 , v2, v3?

OpenStudy (anonymous):

hmm let me think abt this one

OpenStudy (anonymous):

@samjordon is right You can have a look here @imranmeah91 http://tutorial.math.lamar.edu/Classes/LinAlg/OrthonormalBasis.aspx

OpenStudy (anonymous):

neways it wldnt make a diff what v1 v2 or v3 is since they r orthonormal so when u wld find its magnitude it equals one so sqrt(21) is teh answer

OpenStudy (anonymous):

haha still doubting my answer. it seems correct but i am not sure i am gonna go get my linear algebra notes

OpenStudy (amistre64):

y is a dot product so it has one value in the end

OpenStudy (amistre64):

assume y = a.v what does |a.v| represent ?

OpenStudy (anonymous):

magnitude of dot product

OpenStudy (amistre64):

so absolute value of a dot product as well; and there are no values known for the vectors?

OpenStudy (anonymous):

no, they just said the vector sets was orthonormall basis

OpenStudy (amistre64):

the nulspace represents a plane equation

OpenStudy (amistre64):

hmm, id prolly have to try out a few ortho normals and compare the results to see if the differ by a pattern

OpenStudy (anonymous):

thanks to two of you

OpenStudy (amistre64):

100 010 001 is the standard orthnormal

OpenStudy (amistre64):

|-1| = 1 in that case

OpenStudy (anonymous):

just looked up my notes i think i know what to do

OpenStudy (anonymous):

when u have ||cv_1|| c||v_1|| like u pull out the c

OpenStudy (anonymous):

does that make sense amistre?

OpenStudy (amistre64):

1 3 -1 2 -1 4 1 -1 -7 is orthogonal; and if we normal it .... 1/sqrt(6) 2/sqrt(6) 1/sqrt(6) 3/sqrt(11) -1/sqrt(11) -1/sqrt(11) -1\sqrt(66) 4\sqrt(66) -7\sqrt(66) if that dot product is 1 id say the likelihood of my random guess would match up by accident

OpenStudy (amistre64):

the c thing? sure; pull out a scalar

OpenStudy (anonymous):

sooo like wld u say that the number in front of the V_1, V_2 and V_3 are scalar?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

ohhhh so u wld just pull them out i guess the magnitude wld be 8

OpenStudy (anonymous):

oh and c is absolute value when u pull it out it will always be positive that is what i saw

OpenStudy (amistre64):

or just pull out the positive part and let the negative stay in

OpenStudy (anonymous):

ya i guess

OpenStudy (anonymous):

idk like i just took this course like 4 months ago and i am getting sooo confused

OpenStudy (anonymous):

hahah the more i think abt this the more confused i am getting

OpenStudy (amistre64):

as is the dotproduct represents a sphere or ball stuck at the origin on the end of a vector

OpenStudy (amistre64):

cos(t) = a.v/|1||1| is the angle between any 2 vectors so |a.v| </ 1

OpenStudy (anonymous):

alright, I will read it over, thank

OpenStudy (anonymous):

This will help http://www.wolframalpha.com/input/?i=y%3D2+v1-4v2%2Bv3++find+%7C%7Cy%7C%7C

OpenStudy (anonymous):

lol, thanks lily